CYCLOPROPYLACETYLENE STUDIED IN COLD FREE JET EXPANSION, ROOM TEMPERATURE GAS, AND DILUTE SOLUTION: TIER MODEL IVR

\underline{PAM L. CRUM}, GORDON G. BROWN, KEVIN O. DOUGLASS, BRIAN C. DIAN JAMES E. JOHNS, PRADEEP M. NAIR, HYUN S. YOO, AND BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904..

The vibrational dynamics of the first excited state of the acetylenic C-H stretch of cyclopropylacetylene (CPA) in both gasand solution-phase are investigated. FTMW-detected infrared spectroscopy in a free jet expansion are used to investigate the dynamics of ultracold molecular-beam CPA, while ultrafast time-domain transient absorption spectroscopy is used to study the dynamics of CPA in gas and dilute solution at room-temperature. Good agreement among the results measured by different techniques was reported previously and improved time-domain results have recently provided excellent agreement as well as new dynamic information. Both FTMW-detected IR spectroscopy and femtosecond pumpprobe spectroscopy show a very fast (˜ 0.6 ps) initial IVR rate of the first excited state of the acetylenic C-H stretch and oscillations on the timescales of 5 ps and 20 ps, which is evidence of partial population return into the first excited state of the C-H stretch of CPA. When CPA is solvated (0.05M CCl_{4} solution), the IVR rate maintains the room-temperature isolated molecule value. The solvent contribution to the total relaxation rate in solution (k_{T}_{O}_{T} = k_{I}_{V} _{R} + k_{V} _{E}_{R}) is minor and the main effects of the solvent are attributed to pure dephasing effects that destroy the intramolecular coherence.